### v2

# Beam Splitter Physics and Testing Bell Inequalities with Phase Retarders

Abstract
The role of Beam Splitters (BS) is crucial for quantum physics as it reveals the statistical behavior of quantum mechanical particles (quantons) and some of the fundamental quantum phenomena such as quantum superposition and randomness. Here, we investigate the use of BS and Phase Retarders (P) in combination and emphasize the importance of BS-P-BS systems for emerging quantum technologies. We demonstrate that the detection probabilities are equivalent to the spin measurement probabilities in analog setups. Then, by extending the discussion to spatially correlated two-quanton systems, we construct a suitable basis for analysis of other quantum mechanical processes that seem to have common origins. Last, we propose a new experimental setup for testing Bell-CHSH inequalities with spatially correlated two-quanton BS-P-BS systems by introducing topological phases.
preprint: APS/123-QED
Beam splitters (BS) are the most crucial components of theoretical or experimental studies that demonstrate the behaviors of photons as particles and the existence of physical properties like non-locality or indistinguishability. The main purpose of using BS appears to be quite straightforward from the classical perspective; however, its significance in quantum physics is far more profound [1]. Considering the input and output ports of the BS as quantum states, one can simulate quantum transport [2], create entangled quantum states [3], or build a universal quantum computer [4].
On the other hand, special properties emerge when BSs are utilized in combination with phase retarders (P). BS with P systems (BS-P-BS) reveal strong connections between spin measurement and detection of the quantum mechanical particles (quantons). Accordingly, the probabilities of being spin up or down in an arbitrary direction are equivalent to the probabilities of quantons being detected by detectors. Also, the phase of the retarder corresponds to the angle of the direction in which the spin of quantons is measured.
The importance of using BS-P-BS is far more evident in systems consisting of two spatially correlated (or analogously spin entangled) quantons. Such a two-quanton setup can be designed in two different configurations as either quantons produced from the same source reach the BSs after Ps or into the BS-P-BS system in an EPR-Bohm-type setup [5-6]. Since the BS-P-BS structure contains closed trajectories for each quanton, unlike the other configuration, it allows the study of other quantum mechanical/physical effects that seem to have common origins. Hence, the effects of topological phases on spatially correlated systems can be investigated by introducing electromagnetic field sources at the singular region where the quantons are not allowed to move. In this way, the detection probabilities can be compared with the results in a recent study for spin-entangled quantons in the framework of entirely relativistic quantum mechanics [7]. Moreover, at this stage, it is possible to define an observable in terms of the detection probabilities and to find the expectation values and correlation functions. Thus, the EPR-Bohm-type BS-P-BS setup can be used as a new experimental test method for Bell-CHSH inequalities (BI) ensuring a limit expected to be satisfied by deterministic local-hidden-variable-models (LHMV) [8-9].
Figure 1: (a) Two ports symmetric and lossless (50:50) BS. (b) Schematic representation of the BS-P-BS system
The first indication of this was provided in a different sense by Silverman [10]. In this work, Silverman aimed to reveal the amplitudes for joint-detection probabilities of quantons arrival at two detectors without P, and the effects of spin and Aharonov-Bohm (AB) phase on these probabilities by taking into account reflection (r) and transmission (t) coefficients of BS [11]. He proposed various EPR-Bohm-type hybrid setups consisting of a pair of spatially correlated quantons generated from the same source to obtain the correlation functions that can be controlled by the magnetic flux sourced by solenoids placed at singular points in closed trajectories formed by mirrors and spin-selective BS (SBS). The outcomes of Silverman’s work are quite inspiring in many respects. However, the AB phase is a spin-independent effect, and the proposed setup only allows us to find the detection probabilities without any spin dependency [7]. Hence, the relative phase information does not appear in the expectation value expressions.
In this letter, we demonstrate that the joint-detection probabilities are equivalent to the joint-spin measurement probabilities for all kinds of quantons (bosons, fermions) by introducing BS-P-BS systems. Accordingly, we reveal that the angles of the arbitrary directions of the spin measurements can be arranged by the phases of the retarders. Furthermore, we modify [12] with the Aharonov-Casher phase that arises as a result of the motion of the magnetic dipole moment carrier neutrons around the linear charge distributions by illustrating the effects of topological phases on two-quanton correlated systems. Last, we propose a new gedanken experimental setup for testing BI in BS-P-BS systems.
Beam Splitters.—Consider a symmetric and lossless beam splitter (BS) with two ports 00 and $1$ for incoming, and two ports $D_{0}$ and $D_{1}$ for outgoing quantons (See Fig. 1.a). The basis $(1\thinspace,0)^{T}\equiv\left|0\right\rangle$ and $(0\thinspace,1)^{T}\equiv\left|1\right\rangle$ for ports 0 and 1 can be used to describe the state $\left|\psi^{i}_{in}\right\rangle$ of incoming quantons entering such a BS through the port $i$ $(i=0,1)$. Similarly, to represent the output states one can choose the basis $(1\thinspace,0)^{T}$ and $(0\thinspace,1)^{T}$ for $\left|D_{0}\right\rangle$ and $\left|D_{1}\right\rangle$. Hence, the BS performs a linear transformation that converts $\left|\psi^{i}_{in}\right\rangle$ to $\left|\psi_{out}\right\rangle$. By denoting purely real $(r_{i}=R_{i})$ and purely imaginary $(t_{i}=iT_{i})$ as reflection and transmission amplitudes of the BS respectively (Namely, for incoming quantons through port 0 as $r_{0}$ and $t_{0}$, and through port 1 as $r_{1}$ and $t_{1}$), $\left|r_{0}\right|=\left|r_{1}\right|=\left|t_{0}\right|=\left|t_{1}\right|$; $\left|r_{i}\right|^{2}+\left|t_{i}\right|^{2}=1$; $r_{0}^{*}t_{1}+t_{0}^{*}r_{1}=0$ are provided. Therefore the operation of the BS can be given by a unitary matrix [1],
$BS=\left(\begin{array}[]{cc}r_{0}&t_{1}\\ t_{0}&r_{1}\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}1&i\\ i&1\end{array}\right)$ (1)
Thus, in the case of single incoming quanton through port $i$, the output state is $\left|\psi^{0(1)}_{in}\right\rangle\rightarrow\left|\psi_{out}\right\rangle=\frac{1}{\sqrt{2}}\left[\left|D_{0(1)}\right\rangle+i\left|D_{1(0)}\right\rangle\right]$, and the detection probabilities are 1/2 for each detector.
Figure 2: A proposal for the use of BS-P for spatially correlated two-quanton systems generated from the same source
On the other hand, some significant physical characteristics appear when a second BS is added with a phase retarder, and the detection probabilities of quantons become dependent on the phase of the retarder (see Fig. 1.b). Furthermore, this phase corresponds to the angle of the arbitrary direction on which spin measurements are performed to find the probabilities (up or down). According to the existing configuration, wlog retarder operator ${P}$ can be expressed as,
$P=\left(\begin{array}[]{cc}e^{i\vartheta}&0\\ 0&1\end{array}\right)\thinspace\thinspace$ (2)
In the case of single incoming quanton through port $i$, system evolves into the state $\left|\psi^{i}_{in}\right\rangle\rightarrow\left|\psi_{out}\right\rangle={[BS_{2}][P][BS_{1}]}\left|\psi^{i}_{in}\right\rangle$. After $\mathrm{BS_{1}}$ the initial state of the system transforms into $\left|\psi^{i}_{in}\right\rangle\rightarrow\frac{1}{\sqrt{2}}\left[\left|\psi_{0}^{{}^{\prime}}\right\rangle+i\left|\psi_{1}^{{}^{\prime}}\right\rangle\right]$, and gains an additional phase factor via the retarder $\left|\psi_{0}^{{}^{\prime}}\right\rangle\rightarrow e^{i\vartheta}\left|\psi_{0}^{{}^{\prime}}\right\rangle=\frac{e^{i\vartheta}}{\sqrt{2}}\left[i\left|{D_{0}}\right\rangle+\left|{D_{1}}\right\rangle\right]$. Consequently, the final state of the system can be given in a compact representation,
$\left(\begin{array}[]{cc}\left|0\right\rangle\\ \left|1\right\rangle\end{array}\right)\rightarrow\left|\psi_{out}\right\rangle=\frac{ie^{i\frac{\vartheta}{2}}}{\sqrt{2}}\left(\begin{array}[]{cc}-\sin{\frac{\vartheta}{2}}&\cos{\frac{\vartheta}{2}}\\ \cos{\frac{\vartheta}{2}}&\sin{\frac{\vartheta}{2}}\end{array}\right)\left(\begin{array}[]{cc}\left|{D_{0}}\right\rangle\\ \left|{D_{1}}\right\rangle\end{array}\right)$ (3)
The term $(ie^{i\frac{\vartheta}{2}})$ in (3) is the common phase factor and can be neglected for both two equations. These results are valid for all kinds of quantons (bosons, fermions) and just related with the spatial part of the particle’s wavefunctions and one can easily calculate the probabilities using (3).
Now, we extend the problem to investigate possible uses of BS in correlated systems in two scenarios. First, let’s assume that two spatially-correlated (or simply momentum-correlated) quantons produced from the same source, and follow either orange or blue lines as in the Fig. 2. In this scenario, the initial state of the system can be considered as,
$\left|\psi_{in}\right\rangle=\frac{1}{\sqrt{2}}\left[\left|0\right\rangle_{L}\otimes\left|1\right\rangle_{R}-\left|1\right\rangle_{L}\otimes\left|0\right\rangle_{R}\right]$ (4)
and one can use (2) as a retarder operator with the selection of the phase $\vartheta_{L(R)}$ consistent with the path. The correlated quantons reach BSs after passing through reflecting mirrors and retarders on both sides. Thus, the state of the system (4) transformed into the state $\left|\psi_{out}\right\rangle=(BS_{L}\otimes BS_{R})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. It can be written in an explicit form by neglecting the overall phase factor,
$\left|\psi_{out}\right\rangle$ $=$ $\frac{1}{2\sqrt{2}}([\thinspace|{D_{0}}^{\prime}\rangle+i|{D_{1}}^{\prime}\rangle]\otimes[i|{D_{0}}\rangle+|{D_{1}}\rangle]$ (5)
$-e^{i\vartheta}[\thinspace i|{D_{0}}^{\prime}\rangle+|{D_{1}}^{\prime}\rangle]\otimes[\thinspace|{D_{0}}\rangle+i|{D_{1}}\rangle])$
where $\vartheta_{R}-\vartheta_{L}=\vartheta$. At this stage, joint-detection amplitudes for a quanton at e.g. $D_{0}$ and the other one at $D^{\prime}_{0}$ (or $D^{\prime}_{1}$),
$A(D_{0}^{{}^{\prime}},D_{0})\equiv\left\langle{D_{0}^{{}^{\prime}}}{D_{0}}\right|\left.\psi_{out}\right\rangle=\frac{1}{2\sqrt{2}}i(1-e^{i\vartheta})$ (6)
$A(D^{\prime}_{1},D_{0})\equiv\left\langle{D^{\prime}_{1}}{D_{0}}\right|\left.\psi_{out}\right\rangle=-\frac{1}{2\sqrt{2}}(1+e^{i\vartheta})$ (7)
and the associated probabilities are revealed as:
$P(D_{0}^{{}^{\prime}},D_{0})\equiv\left\|A(D_{0}^{{}^{\prime}},D_{0})\right\|^{2}=\frac{1}{2}\cos^{2}{\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)}$ (8)
$P(D_{1}^{{}^{\prime}},D_{0})\equiv\left\|A(D_{1}^{{}^{\prime}},D_{0})\right\|^{2}=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right).$ (9)
These results are related with the joint-spin measurement probabilities obtained in EPR-Bohm type experimental setups using spin-correlated particles [5-6], with a phase shift $\pi/2$ as Degiorgio’s relation due to BS [13]. It is also clear that the probabilities depend on the phase of the retarder.
Second, a new BS-P-BS system for a correlated quanton pair can be proposed to establish a comprehensive analogy between spin measurements and detection processes. Accordingly, spatially-correlated quantons produced in the state (4) are subjected to retarders placed on one of the arms after passing the first BSs. Then, they move through the second BSs and reach the detectors (See Fig 3). In this new configuration, the final state of the system is obtained as $\left|\psi_{out}\right\rangle=(BS_{L_{1}}\otimes BS_{R_{1}})(P_{L}\otimes P_{R})(BS_{L_{2}}\otimes BS_{R_{2}})\left|\psi_{in}\right\rangle$, or clearly,
$|\psi_{out}\rangle=\frac{1}{\sqrt{2}}[(-\sin{\frac{\vartheta_{L}}{2}}|{D_{0}}^{\prime}\rangle+\cos{\frac{\vartheta_{L}}{2}}|{D_{1}}^{\prime}\rangle)\otimes$
$(\cos{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\sin{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)-(\cos{\frac{\vartheta_{L}}{2}}|D_{0}^{\prime}\rangle+\sin{\frac{\vartheta_{L}}{2}}|D_{1}^{\prime}\rangle)$
$\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\thinspace\otimes(-\sin{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\cos{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)]$ (10)
Thus, one can easily find the joint-detection probabilities for quantons at $D_{0}$ and $D_{1}$ (or $D_{1}^{\prime}$):
$P(D_{0}^{{}^{\prime}},D_{0})=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Uparrow_{\tilde{n}})$ (11)
$P(D_{1}^{{}^{\prime}},D_{0})=\frac{1}{2}\cos^{2}{\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)}\equiv P(\alpha\Downarrow_{\tilde{n}},\beta\Uparrow_{\tilde{n}})$ (12)
The above results are well known expressions for probabilities of joint-spin measurements in entangled quantum systems [7]. The phase of the retarder can be thought of as the angle of the arbitrarily chosen direction $\hat{\bm{n}}$ in which the spin is measured in an analog experimental setup. In this regard, (9) clearly indicates the connections between these processes without any Degiorgio phase shift. Accordingly, the probabilities of spin up (or down) in $\hat{\bm{n}}$ direction are equivalent to the probabilities of detected by a detector placed up (or down) for the corresponding phase of the retarder to the angle of n with horizontal.
Figure 3: The use of the BS-P-BS system based on detection probabilities as an analog to spin measurement probabilities for spatially-correlated two-quanton systems and a proposal for a gedanken experimental setup for testing BI.
Moreover, the proposed setup can be used to test BI which provides a limit expected to be satisfied by the deterministic LHVM [5,9]. The violation of BI is meaningful for demonstrating the existence of non-local properties of quantum mechanics. Therefore, in EPR-Bohm-type experimental setups, one tries to find the relevant correlation function S as a combination of expectation value expressions based on different spin-measurement probabilities. In general, to obtain the correlation function S in an entangled two spin-1/2 particles in a singlet state
$\left|\psi_{in}\right\rangle=\frac{1}{\sqrt{2}}\left(\left|\Uparrow_{\bm{\hat{n}}}\right\rangle_{L}\otimes\left|\Downarrow_{\bm{\hat{n}}}\right\rangle_{R}-\left|\Downarrow_{\bm{\hat{n}}}\right\rangle_{L}\otimes\left|\Uparrow_{\bm{\hat{n}}}\right\rangle_{R}\right)$ (13)
one usually starts with decomposing it into the instantaneous eigenstates of the system in any arbitrarily chosen direction $\bm{\hat{n}}$, then observes the joint-spin measurement probabilities of particles in the directions that make different $\alpha$ and $\beta$ angles with $\bm{\hat{n}}$ on both sides. In our configuration, the spin measurement probabilities is replaced by the detection probabilities.
Figure 4: A proposal for testing the possible effects of the AB phase on entangled quantum states.
As an illustration (See Fig 4), joint-probabilities for the particle on the left finding spin up on the direction making an angle $(\vartheta_{L}\equiv\alpha)$ and for the particle on the right finding on the direction making an angle $(\vartheta_{R}\equiv\beta)$ spin up $P\left(\bm{\alpha}\Uparrow_{\bm{\hat{n}}},\bm{\beta}\Uparrow_{\bm{\hat{n}}}\right)$ or down $P\left(\bm{\alpha}\Uparrow_{\bm{\hat{n}}},\bm{\beta}\Downarrow_{\bm{\hat{n}}}\right)$ with the direction $\bm{\hat{n}}$ are exactly same with the joint-spin probabilities $P(D_{0}^{{}^{\prime}},D_{0})$ and $P(D^{\prime}_{1},D_{0})$ found in (9). By introducing the observable $\mathcal{O}$,
$\mathcal{O}^{l}(\vartheta_{L})=P_{D_{0}^{{}^{\prime}}}^{l}(\vartheta_{L})-P_{D_{1}^{{}^{\prime}}}^{l}(\vartheta_{L})$ (14)
and similarly $\mathcal{O}^{r}(\vartheta_{R})$ for the angle $\vartheta_{R}$, one calculate the expectation value of joint measurements,
$E(\vartheta_{L},\vartheta_{R})$ $=$ $\left\langle\psi\left(\tau\right)\right|\mathcal{O}^{l}(\vartheta_{L})\otimes\mathcal{O}^{r}(\vartheta_{R})\left|\psi\left(\tau\right)\right\rangle$ (15)
$=$ $-\cos{(\vartheta_{L}-\vartheta_{R})}$
and the correlation function can be derived to test the BI as,
$S(\vartheta_{L},\vartheta_{R},\vartheta_{L}^{\prime},\vartheta_{R}^{\prime})=\left|E(\vartheta_{L},\vartheta_{R})-E(\vartheta_{L},\vartheta_{L}^{\prime})\right|$
$+\left|E(\vartheta_{R}^{\prime},\vartheta_{R})+E(\vartheta_{R}^{\prime},\vartheta_{L}^{\prime})\right|$ (16)
With the certain chosen angles for the correlation function $S(0,\frac{\pi}{4},\frac{3\pi}{4},\frac{\pi}{2})=2\sqrt{2}>2$, BI is maximally violated by the quantum mechanical expectation values. Historically, violations of BI have been demonstrated in numerous experimental studies using neutral kaons or similar large systems [14-20]. However, the proposed setup is quite simple compared to its counterparts. It is easy to implement in a laboratory environment with all kinds of quantons (bosons or fermions). It has fully closed trajectories on both sides (in contrast to the preceding one), and it enables the investigation of other physical/quantum mechanical systems as topological phases, which is particularly important for experimental studies and emerging quantum technologies. Besides, it is also very noticeable to reach that conclusion based on the probabilities of being detected, without any spin information. In order to use a spin-entangled particle pair instead, it would be sufficient to utilize spin-selective BSs (SBSs) (which reflect all spin down and pass all spin up particles).
Figure 5: (a) Schematic representation of the AC phase in 2+1 dimensions resulting from the motion of magnetic dipole moment carrier chargeless particles in a closed trajectory around a linear charge distribution. (b) A proposal for a gedanken experimental setup to investigate the effects of topological phases on entangled quantum states and to observe the violations of BI using physical quantities in the singular region to test the non-local properties of quantum mechanics.
Testing Bell inequalities with phase retarders.—To provide an illustration of the use of BS detailed in the previous section, we reconsider proposed setups by introducing a topological phase. Hence, we observe the contribution of particle spins to the process and make the function S given by (13) for BI dependent on the topological phase. Topological phases are quantum mechanical phenomena that occur in multiply-connected region under the influence of vector potential or vector potential-like physical quantities, without classical Lorentz forces, independent of the geometry of the space and the speed of the particles [11-12, 21-25]. It is also invariant under spatial translations (traditionally in the z-direction), and essentially arises in two-dimensional space. The first and most well-known example of topological phases is the AB phase, which is based on the control of the interference pattern via an infinitely long and thin solenoid placed beyond the reach of moving electrons in a double-slit experiment. However, the interaction Hamiltonian describing the dynamics of the AB system and the transformed wavefunction are given by,
$\begin{matrix}\Delta H_{AB}=-e\bm{\alpha}\cdot\bm{A}\\ \psi^{\prime}=e^{-ie\Phi}\psi_{0}\end{matrix}$ (17)
where e is the electron charge, $\bm{\alpha}$ is the Pauli matrices, and A is the vector potential in 2+1 dimensions (for detailed information see ref: [7]). The interaction (between the particles and vector potential sourced by the solenoid), and the phase are independent of the spin orientation of the particles as is seen from (14). Thus, the relative phase in spin-entangled systems disappears in the expectation value expressions for the AB case contrarily done in Silverman’s study [10]. The results are reduced to the equations (8) and (9) even in the presence of the solenoid in proposed setups.
Instead, it is appropriate to use the AC phase arising in a stationary reference frame of AB electrons in the presence of unpolarized magnetic dipoles [12]. The AC phase is based on controlling the interference pattern with the linear charge distribution in the singular region in a double-slit experiment with neutrons carrying magnetic dipole moments without the effects of the classical forces. The AC Hamiltonian describing the interaction, and the transformed wave function are as follows,
$\begin{matrix}\Delta H_{AC}=-s\bm{\mu}\cdot\bm{\tilde{E}}\\ \psi^{\prime}=e^{-is\mu\lambda}\psi_{0}\end{matrix}$ (18)
[…]
$\begin{matrix}\left|\Uparrow_{\tilde{n}};t\right\rangle=\frac{1}{\sqrt{2}}(-e^{-i\theta},1)^{T}\\ \\ \left|\Downarrow_{\tilde{n}};t\right\rangle=\frac{1}{\sqrt{2}}(e^{-i\theta},1)^{T}\end{matrix}$ (19)
and the corresponding eigenvalues are $\lambda_{AC_{1,2}}=s\mu\tilde{E}$ with the definitions $\tilde{E}_{x}=\tilde{E}\cos{\theta},\tilde{E}_{y}=\tilde{E}\sin{\theta}$ and $\tilde{E}^{2}=\tilde{E}_{x}^{2}+\tilde{E}_{y}^{2}$, where $\theta$ is the angle between positive $x$ axis and arbitrary $\bm{\hat{\tilde{n}}}$ direction (See Fig. 5.a). Each eigenstate gains a topological AC phase associated with their spin orientation after a complete cycle at $t=\tau$,
$\begin{matrix}\left|\Uparrow_{\tilde{n}};t=0\right\rangle\to\left|\Uparrow_{\tilde{n}};t=\tau\right\rangle=e^{-i\mu\lambda}\left|\Uparrow_{\tilde{n}};t=0\right\rangle&\\ &\\ \left|\Downarrow_{\tilde{n}};t=0\right\rangle\to\left|\Downarrow_{\tilde{n}};t=\tau\right\rangle=e^{i\mu\lambda}\left|\Downarrow_{\tilde{n}};t=0\right\rangle\end{matrix}$ (20)
At this stage, above equations provide a suitable basis for the study of entangled quantum states, and allows the setups proposed in the previous section to be developed with the AC phase.
Now, first let us verify that the effect of the AC phase cannot be observed with a source of the electrical field linear charge distribution placed in the setup given in Figure 2. With the selection of initial state (4) for the spatially-correlated (or spin-entangled) particle pair, the final state of the system is $\left|\psi_{out}\right\rangle^{\prime}=({BS_{L}}\otimes{BS_{R}})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. Here, $T_{L(R)}$ is the quantum topological phase operator,
[…]
$\left|\psi_{out}\right\rangle^{\prime}=e^{i\mu\Delta\lambda}\left|\psi^{(4)}_{out}\right\rangle$ (22)
The probabilities of detection (or spin measurements) for quantons are the same as (7). Hence, the overall phase factor disappears in the expectation value expressions as predicted.
Second, in order to observe such effects, the physical/quantum mechanical processes in Fig. 3 can be improved by adding linear charge distributions (See Fig. 5.b). Similarly, by choosing (4) as the initial state, the output state of the system is $\left|\psi_{out}\right\rangle=({BS_{L}}_{2}\otimes{BS_{R}}_{2})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})({BS_{L}}_{1}\otimes{BS_{R}}_{1})\left|\psi_{in}\right\rangle$. By neglecting the overall phase factor, and taking into consideration (3) and (18), the final state of the system can be rewritten explicitly as ($\lambda=\lambda_{L}-\lambda_{R}$),
$|\psi_{out}\rangle$ $=$ $\frac{1}{\sqrt{2}}[(-\sin{\frac{\vartheta_{L}}{2}}|{D_{0}}^{\prime}\rangle+\cos{\frac{\vartheta_{L}}{2}}|{D_{1}}^{\prime}\rangle)\otimes(\cos{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\sin{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)$ (23)
$-e^{-2i\mu\lambda}(\cos{\frac{\vartheta_{L}}{2}}|D_{0}^{\prime}\rangle+\sin{\frac{\vartheta_{L}}{2}}|D_{1}^{\prime}\rangle)\otimes(-\sin{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\cos{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)]$
The phase factor in (20) appears in the expectation value expressions in contrast to (19), and it allows the correlation function S to be controlled by a physical quantity placed in the singular region without any classical interaction. To demonstrate this, we first find the joint-detection probabilities (or analogously the joint-spin measurement probabilities for arbitrary $\bm{\hat{\tilde{n}}}$ directions) which reveals the statistical behavior of quantons,
$P(D_{0}^{{}^{\prime}},D_{0})\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Uparrow_{\tilde{n}})$
$=\frac{1}{4}[1-\cos{\vartheta_{L}}\cos{\vartheta_{R}}-\sin{\vartheta_{L}}\sin{\vartheta_{R}}\cos(2\mu\lambda)]$
$P(D_{0}^{{}^{\prime}},DD_{1})\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Downarrow_{\tilde{n}})$
$=\frac{1}{4}[1+\cos{\vartheta_{L}}\cos{\vartheta_{R}}+\sin{\vartheta_{L}}\sin{\vartheta_{R}}\cos{(2\mu\lambda)}]$ (24)
Then, by introducing observables $\mathcal{O}^{L(R)}$ as in equation (11), the expectation values are revealed,
$E(\vartheta_{L},\vartheta_{R})=-\cos{\vartheta_{L}}\cos{\vartheta_{R}}-\sin{\vartheta_{L}}\sin{\vartheta_{R}}\cos(2\mu\lambda)$ (25)
Last, by substituting (22) into (13), with the certain chosen angles the correlation function S is obtained as,
$S(0,\frac{\pi}{4},\frac{3\pi}{4},\frac{\pi}{2},\lambda)=\sqrt{2}+\sqrt{2}|\cos(2\mu{\lambda)}|$ (26)
Equation (20) illustrates how the S function depends on the physical quantities in the singular region. Accordingly, for specific angle selections, the S function can be controlled by the topological phases. In the case of a linear charge distribution $\lambda=0$, it reduces to $S=2\sqrt{2}$. Moreover, these are identical to the results of [7] obtained in a spin-correlated AC system and to the results of [Bertlman], where the Berry phase is studied in the case of confined azimuthal angles. The fact that these results are analogous to spin measurements based on detection probabilities shows the importance of BS-P-BS systems and BS applications in emerging quantum technologies.
Conclusion.—The main aim of this study is to analyze the characteristics arising from the use of BS in combination with P. In BS-P-BS systems, we first demonstrate that the detection probabilities of quantons depend on the phase of P, and these probabilities are equivalent to the probabilities of spin measurement in arbitrary directions in analog setups. Here, the phase of P is equivalent to the angle of the direction in which the spin is measured. Then, we propose two possible EPR-Bohm-type experimental setups for developing this use of BS for two-quanton systems.
In the first scenario, spatially-correlated quantons generated from the same source pass directly through P and reach the BSs. The (7) probabilities are related to the joint-spin measurement probabilities in EPR-Bohm type setups (with a $\frac{\pi}{2}$ phase difference as in the diffraction experiments). On the other hand, the effects of the AB phase on the spatially-correlated quanton system were studied by Silverman in a similar proposed setup without P, where the phases are predicted to affect the probabilities of detection (see [10] equations 4-5). However, the paths of the moving quantons do not form a closed trajectory, and it is not a suitable setup for observing the topological phase effects. Moreover, since the AB (and Dual AB) phases are independent of the spin orientations of the incoming quantons, the effects of these phases on spin-entangled states do not arise in the expectation value expressions [26-28]. In conclusion, although the results of Silverman’s work are inspiring, the relevant probabilities cannot be obtained using this setup, and it needs to be modified by the P and spin-dependent AC (or HMW) phase [23-25].
Secondly, a BS-P-BS system is proposed in which two spatially-correlated quantons produced from the same source follow closed trajectories (as in interference experiments). The results demonstrate that the joint-detection probabilities are equal to the probabilities of joint-spin measurements for spin-entangled quantons in EPR-Bohm-type setups. Accordingly, the spin-up and spin-down probabilities for measurements in arbitrary directions in spin-entangled systems are equal to the probabilities of detection by the up-placed detector and the down-placed detector for the determined phase of retarders analogously. These are crucial results and allow us to test the BI. Thus, one can reach the expectation value expressions and the correlation function S, which depends on detection probabilities and the phase of the retarder by introducing an observable. Here, the violation of BI can be followed by controlling P.
Moreover, in contrast to the previous setup, the correlated BS-P-BS system is appropriate for investigating the effects of topological phases. In this regard, the BS-P-BS configuration is reconsidered in the presence of the spin-dependent AC (or HMW) phase. Now, the obtained probabilities depend on the physical quantity in the singular region as well as the phase of the retarder and identical to reference [7], where the effects of topological phases are examined in the framework of entirely relativistic quantum mechanics using spin-entagled systems in 2+1 dimensions. Furthermore, geometric phases that occur in both classical and quantum systems adiabatically evolving through a closed cycle defined in the configuration space of its external parameters are special cases of topological phases [29-31]. Hence, it is necessary to verify the obtained results for the study by Bertlman, where the effects of geometric phases on the entanglement quantum states are investigated [32-34]. Indeed, it is understood that the probability expressions for [32] results will be reduced to expression (21) by confining the azimuthal angles for 2+1 dimensions. This is also a correction to [10] and gives the probability expressions intended to be found. From this point, the intended results can be achieved by choosing the matrix elements,
$BS=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}-\sin{\frac{\vartheta_{L}}{2}}&-i\sin{\frac{\vartheta_{R}}{2}}\\ i\cos{\frac{\vartheta_{L}}{2}}&\cos{\frac{\vartheta_{R}}{2}}\end{array}\right)$ (27)
for r and t coefficients analogous to the phase of the retarder. Although (24) implies that the results can be obtained using SBSs that are not symmetric, this would be unusable in practice. The necessity of choosing a different SBS each time does not satisfy the spin measurement analogy. Last, the proposed setup is utilized as a new testing method for BI. The correlation function becomes dependent only on the physical quantity in the singular region with the phases (angles) that are maximally violated by quantum mechanical expectation values. Consequently, the non-local properties of quantum mechanics can be tested simply using linear charge distribution and magnetic dipole moment carrier chargeless quantons.
The author would like to thank Prof. Ali Ulvi Yilmazer, Prof. Anatoli Polkovnikov, Dr. Melik Emirhan Tunalıoğlu and Dr. Tatsuhiko Ikeda for their valuable contributions. This work is supported by TUBITAK-BIDEB 2219 Research Fellowship.